The life of a compressor manufactured by a company is known to be 200 months on an

average following an exponential distribution. Find the probability that the life of a

compressor of that company is (i) less than 200 months (ii) between 100 months to 25

years?


1
Expert's answer
2021-06-14T15:22:58-0400

Exponential Distribution Probability 



"F(x;\lambda) = \begin{cases}
 0 &\text{if } x<0 \\
 1-e^{-\lambda x} &\text{if } x\geq0
\end{cases}"

Given "\beta=200, \lambda=\dfrac{1}{\beta}=\dfrac{1}{200}"

 i) less than 200 months


"P(X<200)=F(200;\lambda)=1-e^{-{1 \over 200}( 200)}"




"=1-e^{-1}\approx0.63212"



ii) between 100 months to 25 years.



"P(100<X<3000)=F(3000;\lambda)-F(100;\lambda)"




"=1-e^{-{1 \over 200}( 3000)} -(1-e^{-{1 \over 200}( 100)} )"




"=e^{-{1 \over 2}}-e^{-15}\approx0.60653"